OpenStudy (ohmybookness):

Need help with Conic Sections!! Will give medal! Question in comments. Please any help is appreciated.

OpenStudy (ohmybookness):

I solved everything else I just need to solve #6-10. it says to solve it algebraically & to use a graph to demonstrate that my answers are correct. I believe I already have a graph I just need help solving them algebraically.

OpenStudy (ohmybookness):

@ckylee can you help me here?

OpenStudy (ohmybookness):

@ganeshie8 can you help me with these questions?

OpenStudy (anonymous):

For #6, First use elimination to find x and y values \[32x^2 + 9y^2 = 324\]\[3y^2 - x^2 = 3\] \[32x^2 + 9y^2 = 324\]\[-3x^2 + 9y^2 = 9\] \[35x^2 = 315\] \[x^2 = 9\]\[x = \pm3\]\[y= \pm2\] (plug x=3 to any equation to get y) Therefore there are combinations of (3,2), (3, -2), (-3, 2) and (-3,-2).

OpenStudy (anonymous):

I'm not sure about conic sections (ellipses and circles) though. Trigonometry isn't my strongest area :P